the sum of reciprocal of rehman age[in years],3 years ago 5 years from now is 1/3.find his present age.
Answer by dkppathak(439)

the sum of reciprocal of rehman age[in years],3 years ago 5 years from now is 1/3.find his present age.

let present age of rahaman is x yrs
3 yrs ago age of rahman was x-3 yrs
5 yrs from now age of rahman will be x+5 yrs
as per given conditions
1/x-3 +1/x+5 =1/3
by LCM
x+5+x-3/(x+5)(x-3) =1/3
2x+2/x^2 +2x -15 =1/3
by cross multiplication we can find
6x+6 =x^2 +2x-15
x^2+2x -15 -6x -6 =0
x^2 -4x-21=0
(x-7)(x+3) =0
x-7=0 or x=3 =0
x=7 or x=-3
neglecting negative value of age
we can conclude
rahman present age is 7 yrs